OpenStudy (aravindg):

et f(x)=e^x and g(x)=log x find a (f+g)(1) b (fg)(1) c (fog)(1) d (2f)(1)

OpenStudy (anonymous):

(F+G)(1)=e+0=e (FG)(1)=e∗0=0 (F o G)(1)=e0=1 2F(1)=2e

OpenStudy (amistre64):

please do not post your questions inside other peoples questions, that is considered rude and intrusive to say the least

OpenStudy (aravindg):

k